OpenStudy (anonymous):

Find the difference of 9 radical 7 and radical 28.

OpenStudy (lgbasallote):

"radical" 28 = square root of 28 = \(\sqrt {28} = \sqrt{(2)(2)(7)} = \sqrt{(4)(7)}\) that giving you any ideas so far?

OpenStudy (anonymous):

kinda

OpenStudy (lgbasallote):

\(\sqrt {(4)(7)} = 2\sqrt{7}\) and that you can combine with \(9 \sqrt 7\)

OpenStudy (anonymous):

gotcha thanks